# RobustBlack: Challenging Black-Box Adversarial Attacks on State-of-the-Art Defenses

### v1

@@ 5.2 Results
Non-robust target: The results clearly show that using robust surrogate models is detrimental to attack success rates. Across all black-box attacks, non-robust surrogates consistently outperform robust surrogates, with an average improvement of 35.39 percentage points in attack success rates on the vanilla ResNet50 model. For instance, GHOST improves from 17.95% to 49.36%, BASES from 17.94% to 81.08%, and TREMBA from 67.54% to 89.56%. Even UAP that had the lowest success rate improves from 0.73% to 7.11%.
@@
Robust target: By contrast, robust surrogates yield significant improvement in the success rate of black-box attacks when targeting robust models. Notably, attacks using robust surrogates on top-tier robust models (e.g., Liu ConvNeXt-L and Liu Swin-L) yield up to 15 times higher success rates than their non-robust counterparts. For instance, when using the GHOST attack to target the highly robust Swin-L model, the robust surrogate achieves a success rate of 8.16%, compared to just 0.53% with the non-robust surrogate. Similarly, BASES improves its success rate from 1.26% to 12.55% against the Swin-L model. The significant improvement of success rate is consistent across all defense mechanisms, and all attacks, except LGV, and UAP. We hypothesize that the initial training phase of the LGV models, where a high learning rate is used over additional epochs, potentially weakens the model’s robustness during the early sampling phase, which contrasts with GHOST networks that maintain stronger robustness after their creation. This hypothesis is explored in Appendix D.
@@
The results highlight that the choice of surrogate model plays a critical role in the effectiveness of transfer-based black-box attacks, particularly in restricted environments where the robustness of the target is unknown. Robust models, by design, tend to secure blind spot regions in the input space in which vanilla models often fail to predict adversarial examples correctly. As a result, transfer attacks on robust surrogates aim to fool other regions of the input space that are not yet secured by adversarial training, making them better suited for attacking robust models. On the other hand, transfer attacks on vanilla surrogates focuses more on blind spot regions of the input space that are also likely to be mispredicted by other vanilla target models.
@@
Insight 3 The choice of the surrogate model, robust or vanilla, impacts the success of black-box attacks. Robust surrogates are more effective against robust targets, while vanilla surrogates excel against vanilla targets. Attackers can therefore improve attack effectiveness by adaptively selecting surrogates based on the perceived robustness of the target model.
6 Conclusion and Perspectives
@@
This research investigated the effectiveness of black-box adversarial attacks against SoTA and standard robust models using ImageNet, structured around three series of experiments, uncovering key insights.. Our first findings demonstrate that even advanced black-box attacks often struggle against a relatively simple adversarial defense. This suggests that adversarial training can be a highly effective baseline defense for real-world systems accessible via APIs. Consequently, understanding the limitations of current black-box attacks against robust models is crucial for developing more effective attack strategies. Based on our second experiments, we found that defense mechanisms optimized against AutoAttack often generalize effectively to resist black-box attacks, even though performance differences are less pronounced in black-box settings. This finding supports the relevance of AutoAttack as a benchmark for black-box robustness and suggests that current efforts focused on white-box defenses provide a meaningful degree of robustness in black-box scenarios as well – though ensemble defenses may not always translate their AutoAttack robustness to black-box robustness, especially against attacks using similar surrogate models as the ensemble. Finally, we observed that selecting a surrogate with a level of robustness aligned with that of the target can significantly impact the success rate of transfer-based attacks. This finding reveal the possibility for strategic selection of surrogate models when planning black-box attacks. Future work should consider designing black-box attacks that specifically target SOTA robust models to better reflect real-world defensive capabilities. Additionally, further exploration of why SoTA defenses optimized for AutoAttack are also relevant for black-box robustness could yield valuable insights into cross-attack resilience mechanisms. Expanding the study of transfer-based attacks to consider varying levels of robustness between surrogate and target models would also help to deepen our understanding of transferability dynamics. Finally, while this study focused on image classification tasks, future research could extend the analysis to other tasks, such as object detection or segmentation, to understand how black-box robustness may vary across applications. In conclusion, this study demonstrates that SoTA defenses are notably resilient against black-box attacks, underscoring the importance of developing more targeted attack strategies to effectively challenge these modern robust models. As the field advances, these insights can guide both the creation of more sophisticated black-box attacks and the design of improved defense mechanisms that support the deployment of robust AI systems across a wide range of applications.
Acknowledgment
This research is supported by the Luxembourg National Research Fund under the AFR grant. We extend our gratitude to Martin Gubri (ParameterLabs) for his invaluable support with the LGV implementation.
Appendix
Appendix A Hyperparameters of Transfer Attacks
Tables 5 and 6 provide a comprehensive overview of the hyperparameters used for the ten transfer-based adversarial attacks in our experiments. They categorize the key hyperparameters into two groups: common hyperparameters shared across most attacks and attack-specific hyperparameters that define the unique characteristics of each method. Most of the attack hyperparameters were kept at their default values, as specified in the repositories or libraries that provide the implementations of adversarial attacks (Kim, 2020; Woo, 2024).
A.1 Common Hyperparameters
Table 5 provides the default values and explanations for these common hyperparameters, such as the step size (Alpha), momentum factor (Decay), and number of iterations (Steps), which are essential for iterative attacks as they control the perturbation magnitude and the optimization process.
Table 5: Common hyperparameters for transfer-based attacks.
PARAMETER DEFAULT VALUE EXPLANATION APPLIES TO
Alpha 2/255 Step size or update rate for perturbation. All attacks except UAP
Decay 1.0 Momentum factor. All attacks except UAP
Steps 10 Number of iterations. All attacks except UAP
Epsilon (eps) 4/255 Maximum perturbation for adversarial example. All attacks
A.2 Attack-Specific Hyperparameters
In addition to the shared hyperparameters, each attack incorporates its own unique hyperparameters tailored to its specific mechanisms. Table 6 summarizes these attack-specific hyperparameters, listing the default values, explanations, and the corresponding attacks to which they apply.
Table 6: Attack-specific hyperparameters for transfer-based attacks.
PARAMETER DEFAULT VALUE EXPLANATION APPLIES TO
Resize Rate 0.9 Rate at which the input images are resized. DIFGSM, TIFGSM
Diversity Prob 0.5 Probability to apply input diversity. DIFGSM, TIFGSM
Kernel Name Gaussian Type of kernel used. TIFGSM
Kernel Length 15 Length of the kernel. TIFGSM
Sigma (nsig) 3 Radius of the Gaussian kernel. TIFGSM
N 5 Number of samples in the neighborhood. VMI, VNI
Beta 1.5 Upper bound for neighborhood. VMI, VNI
Prior Type no_data Type of prior used for attack optimization. UAP
Patience Interval 5 Number of iterations to wait before checking convergence. UAP
Gamma 0.5 Decay factor for gradients from skip connections. SGM
LGV Epochs 5 Number of epochs for training the LGV models. LGV
LGV Models Epoch 2 Number of models used per epoch. LGV
LGV Learning Rate 0.05 Learning rate for the LGV training phase. LGV
Randomized Modulating Scalar 0.22 Drawn from the uniform distribution 1-scalar, 1+scalar. GHOST
Portion 0.2 Portion for the mixed image. Admix
Size 3 Number of randomly sampled images. Admix
Appendix B Hyperparameters of Query Attacks
Tables 7 and 8 provides an overview of the hyperparameters used in BASES and TREMBA, respectively. Most of these hyperparameters were kept at their default values, as specified in their repositories, since our experiments aimed to evaluate their performance under standard configurations. For BASES, key hyperparameters include the ensemble size (N_wb), the type of fusion for surrogate models (Fuse), and the number of weight updates (Iterw). Similarly, TREMBA, which operates in two phases (training and attacking), relies on hyperparameters such as the number of training epochs (Epochs), the number of attack iterations (Num_Iters), the number of samples for Natural Evolution Strategy (NES), and latent space parameters like the Gaussian noise standard deviation (Sigma).
Table 7: Hyperparameters for the BASES attack.
PARAMETER DEFAULT VALUE EXPLANATION
Epsilon (eps) 4 Perturbation bound magnitude (0 to 255 pixel range)).
N_wb 10 Number of models in the ensemble.
Bound linf Perturbation bound norm type.
Iters 10 Number of inner iterations.
Fuse loss Fuse method, e.g., loss or logit.
Loss_name CW Loss function used for optimization.
Algo PGD White-box algorithm used for the perturbation machine
X 3 Factor to scale the step size alpha.
Learning Rate (lr) 0.005 Learning rate for weight updates.
Iterw 20 Number of iterations to update weights.
Table 8: Hyperparameters for the TREMBA attack.
PARAMETER DEFAULT VALUE EXPLANATION
Epsilon (eps) 8/255 (Training Phase), 4/255 (Attacking Phase) Perturbation bound magnitude.
Learning Rate (G) 0.01 (Training Phase), 5.0 (Attacking Phase) for gradient updates, for latent updates
Momentum 0.9 (Training Phase), 0.0 (Attacking Phase) for SGD, for attack on the embedding space.
Num_images 49k (Training Phase) Size of the training set
Epochs 500 Number of epochs for Training.
Schedule 10 Epochs between Training learning rate decay.
Gamma 0.5 Training Learning rate decay factor.
Margin 200.0 (Training Phase), 5.0 (Attacking Phase) Margin for the loss function.
Sample_Size 20 Number of samples for NES
Num_Iters 200 Number of iterations for NES.
Sigma 1.0 Standard deviation of the Gaussian noise for NES.
Learning Rate Min 0.1 Minimum learning rate in the Attacking Phase.
Learning Rate Decay 2.0 Learning rate decay factor in the Attacking Phase.
Plateau_Length 20 Number of iterations to monitor the objective function
Plateau_Overhead 0.3 Allowed loss threshold to change before the decay
Appendix C Increasing the Budget of Black-Box Attacks
We quadrupled (multiplied by 4) the budget of the best black-box attacks in section 3 (BASES and TREMBA) against the robust ResNet50 model, to reassure that black-box attacks reduced effectiveness (compared to their effectiveness against the vanilla ResNet50 model) is not due to a limited budget. For BASES, we increased the number of inner iterations of PGD within the perturbation machine (Iters) to 40. For TREMBA, we increased the Number of iterations for NES (Num_Iters) to 800.
Figure 5 presents the success rates of BASES and TREMBA attacks against the vanilla ResNet-50 model with standard attack budget, and against the adversarially trained ResNet-50 model under both standard and quadrupled attack budgets. Both attacks underperform against the adversarially trained ResNet50 under quadrupled budgets compared to the vanilla model with standard budgets. For instance, BASES achieves 2.56% ASR on the robust ResNet50 Budget (x4) versus 81.08% on the vanilla ResNet50. Similarly, TREMBA’s ASR is at 5.24% on the robust ResNet50 Budget (x4) compared to 89.56% on the vanilla ResNet50. These findings confirm that our earlier results in section 3 were not a consequence of budget limitations.

### v2

@@ V-B Results
Non-robust target: The results clearly show that the use of robust surrogate models is detrimental to the success rates of the attack. Across all black-box attacks, non-robust surrogates consistently outperform robust surrogates, with an average improvement of 35.19 percentage points in attack success rates on the vanilla ResNet50 model. For instance, GHOST improves from 17.95% to 49.36%, BASES from 17.94% to 81.08%, and TREMBA from 67.54% to 89.56%. Even UAP that had the lowest success rate improves from 0.73% to 7.11%. Notably, attacks that leveraged a training phase before the attacking phase (LGV and TREMBA) exhibited a relatively smaller improvement gap compared to other attacks as robust surrogates were already performing well. Specifically, LGV improved from 81.34% to 87.39%, while TREMBA increased from 67.54% to 89.56%.
@@
Robust target: By contrast, robust surrogates yield significant improvement in the success rate of black-box attacks when targeting robust models, resulting in an average improvement of 6.42 percentage points in ASR across attacks and target models. Notably, attacks using robust surrogates on top-tier robust models (e.g., Liu ConvNeXt-L and Liu Swin-L) yield up to 15 times higher success rates than their non-robust counterparts. For instance, when using the GHOST attack to target the highly robust Swin-L model, the robust surrogate achieves a success rate of 8.16%, compared to just 0.53% with the non-robust surrogate. Similarly, BASES improves its success rate from 1.26% to 12.55% against the Swin-L model. An improvement of more than or equal to 2 percentage points in ASR is consistent across all defense mechanisms and all attacks, except LGV and UAP. We hypothesize that the initial training phase of the LGV models, where a high learning rate is used over additional epochs, weakens the model’s robustness during the early sampling phase, which contrasts with GHOST networks that maintain stronger robustness after their creation. We explore this hypothesis in Appendix B-A.
@@
The results highlight that the choice of surrogate model plays a critical role in the effectiveness of transfer-based black-box attacks, particularly in restricted environments where the robustness of the target is unknown. Robust models, by design, tend to secure blind spot regions in the input space in which vanilla models often fail to predict adversarial examples correctly. As a result, transfer attacks on robust surrogates aim to fool other regions of the input space that are not yet secured by adversarial training, making them better suited for attacking robust models. On the other hand, transfer attacks on vanilla surrogates focus more on blind spots of the input space that are also likely to be mispredicted by other vanilla target models.
@@
Insight 3 The choice of the surrogate model, robust or vanilla, affects the success of black-box attacks. Robust surrogates are more effective against robust targets, while vanilla surrogates excel against vanilla targets. Attackers can therefore improve attack effectiveness by adaptively selecting surrogates based on the perceived robustness of the target model.
VI Discussion
VI-A Impact of defense mechanisms on different attack strategies
Against advanced defenses, all attack mechanisms are disrupted both with $\epsilon=4/255$ and with $\epsilon=16/255$ (9), meanwhile some attacks mechanism remain effective under standard adversarial training (SAT): Compared to vanilla models, SAT is sufficient against all our baseline attacks under low $\epsilon=4/255$ budgets. Under higher $\epsilon=16/255$, only attacks that leverage the decision boundary similarity mechanism (Signflip/Square) and the knowledge transfer hypothesis (TREMBA) achieve over 20% success rate.
We explain this observation by the lower impact of SAT on these two mechanisms compared to the other mechanisms. Indeed, SAT alone is not sufficient to disrupt the decision boundaries enough when comparing two models with the same architectures, and the same training data (same ”knowledge”). The data augmentations in defenses A and the ensembles in defenses B, however, disrupt TREMBA, Square and Signflip by shifting the knowledge and the boundaries. Attacks like SQUARE, which rely on querying the boundary to estimate its position, struggle because the boundary geometry of defenses A becomes fundamentally different from the surrogate’s.
Meanwhile, the gradient similarity attacks like MI-FGSM and VMI-FGSM rely on the assumption that the gradient direction is stable and transferable, which is disrupted by SAT (by smoothing the loss landscape). Using extensive data augmentation and model scaling (defenses A), the boundaries, the knowledge and gradients are ”decorrelated” compared to the source model. These defenses also disrupt shared space attacks (DI, GHOST, LGV, UAP) that exploit large, contiguous volumes of empty space in the data manifold to generate adversarial - the large-scale data augmentation fill these spaces, and prevent UAP from finding universal directions of attacks.
Interaction-based attacks (TI-FGSM) and non-robust feature attacks (ADMIX) try to find shortcuts, using respectively low-order (pixel-local) interactions and non-robust features. Our results confirm that these shortcuts are fundamentally disrupted but all adversarial training defenses, making them unreliable for transfer attacks.
VI-B Impact of Adaptive Attacks.
Our third insight confirms that implementing adaptive attacks by including a robust surrogate when crafting the adversarial examples may not be a suitable approach. Although it improves the effectiveness of the attacks (up to 6.42%), the final success rate achieves at best 12.5% using adaptive BASES against the strongest defense (Liu Swin-L). Meanwhile, adapting the attacks using a surrogate with a specific defense decreases the effectiveness of the attack against vanilla models. Therefore the adaptation needed in the black-box setting should be carefully crafted for each target.
To better understand this phenomenon, we measure the gradient similarity across 1,000 ImageNet samples. For each image-label pair, we calculate the gradient of the cross-entropy loss with respect to the input image for both surrogate and target models. We report the cosine similarity between the gradients. This process was repeated for all surrogate-target model pairs, and the similarities were averaged across all samples to produce a similarity matrix in Fig. 6. Each cell in the matrix represents the average cosine similarity between a surrogate model (rows) and a target model (columns), with values ranging from 0 (orthogonal gradients) to 1 (perfectly aligned gradients). Our results demonstrate that robust targets have a higher gradient alignment with robust surrogates than with vanilla surrogates, even when robust surrogates are from very different architectures and sizes.
Figure 6: Gradient similarity between surrogates and targets over 1000 samples.
VI-C Transferability Bounds Under Mild Robustness Mechanisms
Our study focused on SoTA defenses, that have all been designed against strong attacks ($\epsilon=8/255$ in Robustbench). We did not explore the impact of transferability when the models are midly robustified. However, the work promoting mild robustification [82, 83] was justified only by the tradeoff between robustness and accuracy that was ovserved then. Today, however, very large models and extensive data augmentation allows to robustify the models with agressive attacks and strong mechanisms without degradation of their clean performance (quite the opposite [73]).
VII Limitations and Perspectives
While our work investigated a large set of black-box attacks and robust models, we mostly focused on extremely robust models. [84] for instance focused on mildly perturbed examples and uncovered their interactions with simple adversarial training. Expanding our study of transfer-based attacks to consider varying levels of attack strength and robustness levels between surrogate and target models would also help to deepen our understanding of transferability dynamics. In particular, considering randomness in the pre-processing, and defenses including obfuscation and randomness.
In addition, while this study focused on image classification tasks, future research could extend the analysis to other tasks, such as object detection or segmentation, to understand how black-box robustness may vary across applications. These regression tasks raise a new challenge of defining acceptable robustness thresholds, given that the output is not binary. Although there is an established benchmark for adversarial robustness for classification tasks, there is no such benchmark for segmentation tasks.
Another limitation is that while we included large transformers models, we did not explore the robustness of foundations models such as GPT [57], LLaMA [67], CLIP [56], and SAM [39], or their variants. Research on robustification of these models is still in its infancy for a benchmark.
VIII Conclusion
@@
This research investigated the effectiveness of black-box adversarial attacks against SoTA and standard robust models using ImageNet, structured around three series of experiments, uncovering key insights.. Our first findings demonstrate that even advanced black-box attacks often struggle against a relatively simple adversarial defense. This suggests that adversarial training can be a highly effective baseline defense for real-world systems accessible via APIs. Consequently, understanding the limitations of current black-box attacks against robust models is crucial for developing more effective attack strategies. Based on our second experiment, we found that defense mechanisms optimized against AutoAttack often generalize effectively to resist black-box attacks, even though performance differences are less pronounced in black-box settings. This finding supports the relevance of AutoAttack as a benchmark for black-box robustness and suggests that current efforts focused on white-box defenses provide a meaningful degree of robustness in black-box scenarios as well – though ensemble defenses may not always translate their AutoAttack robustness to black-box robustness, especially against attacks using similar surrogate models as the ensemble. Finally, we observed that selecting a surrogate with a level of robustness aligned with that of the target can significantly impact the success rate of transfer-based attacks. This finding reveals the possibility for strategic selection of surrogate models when planning black-box attacks. Future attacks should consider designing black-box attacks that specifically target SOTA robust models to better reflect real-world defensive capabilities. Additionally, further exploration of why SoTA defenses optimized for AutoAttack are also relevant for black-box robustness could yield valuable insights into cross-attack resilience mechanisms. In conclusion, our study demonstrates that SoTA defenses are notably resilient against black-box attacks, underscoring the importance of developing more targeted attack strategies to effectively challenge these modern robust models. As the field advances, these insights can guide both the creation of more sophisticated black-box attacks and the design of improved defense mechanisms that support the deployment of robust AI systems across a wide range of applications.
LLM usage considerations
LLMs were used for editorial purposes in this manuscript, and all outputs were inspected by the authors to ensure accuracy and originality. Similarly, LLMs were used to assist in the process of writing code, and searching for relevant publications and literature reviews. All code and publications were manually verified and tested to ensure coherence in the results.
Acknowledgements
This research was funded in whole, or in part, by the Luxembourg National Research Fund (FNR), grant reference 18886361. Dr. Ghamizi is supported by the Luxembourg National Research Fund (FNR) CORE C24/IS/18942843.
Appendix A Computational Budget of Attacks
Table XIII details the computational budgets of all studied black-box attacks, quantified as the number of model calls required during the attacking phase per entry.
TABLE XIII: Computational budgets of black-box attacks based on model calls
Attack Algorithm Number of Model Calls Explanation
MIFGSM Iterations (e.g., 10) MIFGSM iteratively refines the adversarial image over a number of steps,
each requiring a model call
DIFGSM Iterations (e.g., 10) DIFGSM performs iterative updates with input diversity applied at each step,
requiring a model call per step
TIFGSM Iterations (e.g., 10) TIFGSM performs iterative updates with translation-invariant gradient smoothing
at each step, requiring a model call per step
VMIFGSM Iterations * (N + 1) (e.g., 60) VMIFGSM iterates Steps times. In each iteration, there’s one model call for the adversarial
image and N model calls for the neighboring images to calculate the gradient variance
VNIFGSM Iterations * (N + 1) (e.g., 60) VNIFGSM, similar to VMIFGSM. In each iteration, there is one model call for the Nesterov and
N model calls for the neighboring images to calculate the gradient variance
GHOST Iterations (e.g., 10) GHOST applies Gaussian noise to the skip connection during each iteration
(one Ghost model per iteration), requiring one model call per step
LGV Iterations (e.g., 10) LGV uses an ensemble of models (lgv_nb_models_epoch * lgv_epochs models),
with one model call at each step for each model in the ensemble
Admix Iterations * Scale * Size (e.g., 150) In each step, Admix generates Size admixed images, scales them Scale times,
and then calculates the loss
UAP Max Iterations + Evaluation Calls The primary loop runs for max_iter iterations. Additionally, a function is
(preparing the perturbation), then 1 called periodically to evaluate the current perturbation
TREMBA (num_iters + 1) * sample_size (e.g., 4000) Tremba iteratively refines the adversarial perturbation within the latent space of a generator network.
Each iteration involves generating sample_size perturbations and evaluating their losses
BASES iterw * n_iters * n_wb (e.g., 2,000) BASES uses a surrogate ensemble containing n_wb models and performs an attack
for n_iters iterations during each query (iterw)
A-A Increasing the Computation Budget of Black-Box Attacks
We are keeping the perturbation budget $\epsilon=4/255$, and quadrupled (multiplied by 4) the budget of the best black-box attacks in section III (BASES and TREMBA) against the robust ResNet50 model, to reassure that black-box attacks reduced effectiveness (compared to their effectiveness against the vanilla ResNet50 model) is not due to a limited budget. For BASES, we increased the number of inner iterations of PGD within the perturbation machine (Iters) to 40. For TREMBA, we increased the Number of iterations for NES (Num_Iters) to 800.
Figure 7 presents the success rates of BASES and TREMBA attacks against the vanilla ResNet-50 model with standard attack budget, and against the adversarially trained ResNet-50 model under both standard and quadrupled attack budgets. Both attacks underperform against the adversarially trained ResNet50 under quadrupled budgets compared to the vanilla model with standard budgets. For instance, BASES achieves 2.56% ASR on the robust ResNet50 Budget (x4) versus 81.08% on the vanilla ResNet50. Similarly, TREMBA’s ASR is at 5.24% on the robust ResNet50 Budget (x4) compared to 89.56% on the vanilla ResNet50. These findings confirm that our earlier results in section III were not a consequence of budget limitations.